# Thermodynamics and dynamics of coupled complex SYK models

### v3

@@ VI.3 Comparing with the Hawking-Page transition
Before we end this section, we would like to comment on Fig. 7 which we reproduce from [30, 31]. The analysis was reproduced and generalized to the charged case in [7] where the authors called the temperature of transition as the “Zorro’s” temperature when $Q_{\text{BH}}=0$ on the coexistence line. As highlighted in appendix E, we show that this temperature is indeed the same as Hawking-Page temperature and the horizontal-axis of Fig. 13 in [7] (as reproduced with more details here in Fig. 7 taken from [30], [31]) is the standard Hawking-Page first order phase transition from a maximally chaotic black hole to non-chaotic free radiation (recall this is true only when $Q_{\text{BH}}=0$) as one cools the temperature down to the Hawking-Page temperature $T_{\text{HP}}$ which is referred to as the “Zorro’s” temperature in [7]. We remind the readers again that the “very” low-temperature regime in the phase diagram is based on the assumption that the only $g(t)$ contributes at leading order at the additional rescaling of temperature $T=\overline{T}q^{-2}$ where we write the Green’s function as ${\cal G}(t)\propto e^{g(t)/q+g_{2}(t)/q^{2}+{\cal O}(1/q^{3})}$. See the text below eq. (81) where this point is discussed in detail. That is why we have provided these preliminary calculations in the conclusion & outlook section of our manuscript (i.e., sections VI.2 and VI.3) where we provide a pathway to make these connections robust as the future work.
@@ VI.4 Outlook
or does it extend beyond this, or is it just a coincidence that we found for our coupled case? Given that the kinetic SYK model is integrable, one would imagine this to change upon adding a quadratic term to the above Hamiltonian. This term would dominate in the low-temperature regime, which would at least mean that the model is unlikely to have a maximally chaotic to maximally chaotic second order phase transition, as observed in this work. But without any kinetic term, the future work can be to find if at large-$q$ and low-temperature limits, all coupled SYK systems as in eq. (82) belong to the same universality class as Landau-Ginzburg and if they are also maximally chaotic. This naturally seems plausible from the renormalization group flow arguments where at lower temperatures, the interaction with less fermions will dominate the Hamiltonian but a full analysis for the aforementioned Hamiltonian is yet to be done (see [32] for the Majorana SYK case whose generalization to the complex case will serve as an interesting inquiry).
@@ Appendix A Equivalence of our model with a coupled SYK chain at equilibrium
Therefore from a thermodynamic point of view, our complex SYK dot in eq. (1) is the same as an equilibrium chain of complex SYK dots with nearest-neighbor hopping with uniform, site-independent coupling. This is the justification for using the word “coupled” in the title of this work.
@@ Appendix D Chaos theory of the SYK model
In chaos theory, we wish to analyze the evolution of the systems operator $W(t)$ after a small perturbation $V(0)$ in the initial conditions. The correlation function describing this evolution is the out-of-time-order correlator (OTOC), generally written as
$c(t)=\expectationvalue{[W(t),V(0)]^{2}}_{\beta},$ (114)
[…]
@@
[…]
@@
Figure 9: Equation of state for charged AdS black hole where $T_{H}$ is the Hawking temperature, $Q_{\text{BH}}>0$ is the total charge of the black hole and $T_{\text{Z}}$ is the minimum temperature of the larger black hole. The green dashed lines are the unstable regimes with negative specific heat. The two stable phases are marked with blue and red. Color coding is the same as in Figs. 1, 2 and 5.

### v4

@@ VI.3 Comparing with the Hawking-Page transition
Before we end this section, we would like to comment on Fig. 8 which we reproduce from [31, 32]. The analysis was reproduced and generalized to the charged case in [7] where the authors called the temperature of transition as the “Zorro’s” temperature when $Q_{\text{BH}}=0$ on the coexistence line. As highlighted in appendix E, we show that this temperature is indeed the same as Hawking-Page temperature and the horizontal-axis of Fig. 13 in [7] (as reproduced with more details here in Fig. 8 taken from [31], [32]) is the standard Hawking-Page first order phase transition from a maximally chaotic black hole to non-chaotic free radiation (recall this is true only when $Q_{\text{BH}}=0$) as one cools the temperature down to the Hawking-Page temperature $T_{\text{HP}}$ which is referred to as the “Zorro’s” temperature in [7]. We remind the readers again that the “very” low-temperature regime in the phase diagram is based on the assumption that only $g(t)$ contributes at leading order at the additional rescaling of temperature $T=\overline{T}q^{-2}$ where we write the Green’s function as ${\cal G}(t)\propto e^{g(t)/q+g_{2}(t)/q^{2}+{\cal O}(1/q^{3})}$. See the text below Eq. (86) where this point is discussed in detail. That is why we have provided these preliminary calculations in the conclusion & outlook section of our manuscript (i.e., sections VI.2 and VI.3) where we provide a pathway to make these connections robust as the future work.
@@ VI.4 Outlook
or does it extend beyond this, or is it just a coincidence that we found for our coupled case? Given that the kinetic SYK model is integrable, one would imagine this to change upon adding a quadratic term to the above Hamiltonian. This term would dominate in the low-temperature regime, which would at least mean that the model is unlikely to have a maximally chaotic to maximally chaotic second order phase transition, as observed in this work. But without any kinetic term, the future work can be to find if at large-$q$ and low-temperature limits, all coupled SYK systems as in Eq. (87) belong to the same universality class as Landau-Ginzburg and if they are also maximally chaotic. This naturally seems plausible from the renormalization group flow arguments were at lower temperatures, the interaction with fewer fermions will dominate the Hamiltonian, but a full analysis for the aforementioned Hamiltonian is yet to be done (see [33] for the Majorana SYK case whose generalization to the complex case will serve as an interesting inquiry).
@@ Appendix A Equivalence of our model with a coupled SYK chain at equilibrium
Therefore, from a thermodynamic point of view, our complex SYK dot in Eq. (1) is the same as an equilibrium chain of complex SYK dots with nearest-neighbor hopping with uniform, site-independent coupling. This is the justification for using the word “coupled” in the title of this work.
@@ Appendix D Chaos theory of the SYK model
In chaos theory, we wish to analyze the evolution of the system’s operator $W(t)$ after a small perturbation $V(0)$ in the initial conditions. The correlation function describing this evolution is the out-of-time-order correlator (OTOC), generally written as
$c(t)=\langle[W(t),V(0)]^{2}\rangle_{\beta},$ (119)
where $\langle\dots\rangle_{\beta}$ is the thermal expectation value, which can be written as $\langle{\cal O}\rangle_{\beta}={\operatorname{Tr}}[\rho(\beta){\cal O}]=\frac{1}{Z}{\operatorname{Tr}}[e^{\beta H}{\cal O}]$. It has been noticed that the correlator grows exponentially with some growth rate $\lambda_{L}$ [2],
@@
$\mathcal{F}(t_{1},t_{2})={\operatorname{Tr}}\left[y\;W(t)\;y\;V(0)\;y\;W(t)\;y\;V(0)\right],\quad y\equiv\rho(\beta)^{1/4}$ (121)
@@
Figure 10: Equation of state for charged AdS black hole where $T_{H}$ is the Hawking temperature, $Q_{\text{BH}}>0$ is the total charge of the black hole and $T_{\text{Z}}$ is the minimum temperature of the larger black hole. The green dashed lines are the unstable regimes with negative specific heat. The two stable phases are marked with blue and red. Color coding is the same as in Figs. 1, 2 and 6.